# Learning Domain-Invariant Features for Out-of-Context News Detection

### v1

@@ 6.3. Visualization
To show that ConDA-TTA effectively learns the domain-invariant feature, we adopt TSNE (Van der Maaten and Hinton, 2008) to visualize the multimodal embeddings x (also the input embedding to ConDA-TTA) and the learned domain-invariant feature z projected to the 2-d space (Figure 3).
@@
Figure 3. TSNE visualization of the multimodal feature x and the learned domain-invariant feature z under Cv, Cl $\rightarrow$ M.
@@
Here we illustrate the TSNE visualization of setting Cv, Cl $\rightarrow$ M. We can tell from the figure that feature z from different topics blend better than x, where we see clearer boundaries between different topics. This suggests that our model has removed domain-specific features from x to some extent. This further shows that our approach is able to effectively learn domain-invariant feature, and transfers out-of-context news detection knowledge to the target domain (Table 1).
6.4. Sensitivity Analysis
The contrastive loss and the MMD are two crucial components in our approach. This section briefly studies how their weights, $\lambda_{ctr}$ and $\lambda_{MMD}$, in the loss function affect the performance of ConDA-TTA.
Figure 4. ConDA-TTA’s performance (in Acc) with different $\lambda_{MMD}$ and $\lambda_{ctr}$ values. The legend denotes the target domain.
Figure 4 illustrates ConDA-TTA’s performance with $\lambda_{MMD}=\{0.5,1,2,5\}$ and $\lambda_{ctr}=\{0.1,0.5,1,2\}$. When setting $\lambda_{MMD}$ to 1, our model achieves the best performance in 4 (M, Cl, B, U) out of 7 domain adaptation settings, and achieves close to best performance in the remaining 3 settings. Overall, $\lambda_{MMD}$ being too high causes the accuracy to drop a little. This might be because the model has put comparatively less emphasis on learning the detection of out-of-context news (i.e. the cross-entropy loss in Eq. (9) receives less attention). For $\lambda_{ctr}$, we find that setting it to 0.5 achieves the best performance in 5 settings. When the value goes up, the model performance tends to decrease gradually. We conjecture this is because the similarity (Section 4.2) among original news becomes smaller, making it slightly difficult to find the decision boundary.
7. Conclusion
@@
This paper has proposed ConDA-TTA, a domain adaptive out-of-context news detection model that can adapt to both unlabeled news topics and news agencies. ConDA-TTA first encodes the news using the BLIP-2 multimodal feature encoder and then adopts contrastive learning and MMD to learn the domain-invariant features. During test-time, we further incorporate target domain statistics into the classifier of ConDA-TTA, in order to better adapt the model to the target domain. Experimental results show the efficacy of our approach. These advantages enable ConDA-TTA to transfer the learned out-of-context news detection knowledge to emerging events or news platforms, lessening the burden of online news content moderators.

### v2

@@ 6.3 Visualization
To show that ConDA-TTA effectively learns the domain-invariant feature, we adopt TSNE Van der Maaten and Hinton (2008) to visualize the multimodal feature x and the learned domain-invariant feature z projected to the 2-d space (Figure 3).
@@
Figure 3: TSNE visualization of the multimodal feature x and the learned domain-invariant feature z under Cv, Cl $\rightarrow$ M.
@@
We can tell from the figure that feature z of different topics blend better than x. Additionally, in Table 3, we compute the variance of x and z after projecting and normalizing them onto 1-d space. The decrease of variance and the TSNE visualization suggest that our model has removed domain-specific features from x to some extent. This further shows that our approach effectively learns domain-invariant features. The parameter sensitivity analysis can be found in Appendix A.3.
Table 3: Variance of x and z on different domains.
Feature Var(M) Var(Cl) Var(Cv)
x 0.0623 0.0738 0.0780
z 0.0326 0.0652 0.0724
7 Conclusion
@@
This paper has proposed ConDA-TTA, a domain adaptive out-of-context news detection model that can adapt to both unlabeled news topics and news agencies. ConDA-TTA first encodes the news using the BLIP-2 multimodal feature encoder and then adopts contrastive learning and MMD to learn domain-invariant features. During test-time, it further incorporates target domain statistics into the classifier. Experimental results show the effectiveness and superiority of our approach.
Limitations
Despite the effectiveness of our proposed ConDA-TTA model, there are several limitations. First, in this work, we focus on two domain types: news topics and news agencies. We do not consider other potential types of domain, such as news from different regions, languages, or cultural backgrounds. Second, we observe that under some domain adaptation settings, all experimented models exhibit negative transfer. In this paper, we present analysis based on the findings in the underlying dataset paper. We will leave a deeper analysis into our future work.
Appendix A Appendix
A.1 Implementation Details
Augmentation: The performance of contrastive learning is largely influenced by the transformation used to generate positive samples (Tian et al., 2020). Following the data augmentation selection in  Bhattacharjee et al. (2023), we experiment with a set of data augmentation techniques and select the one with the best performance on the test set as the final strategy. In view of the large size of the training set (Section 5.1), we randomly sample 1.5% data items from the training set as a toy training set, roughly the same size as the test set. We do this in order to expedite the selection of positive sample generation strategy. We experiment with a set of augmentations: text-level: {synonym replacement, random swap, random crop}, image-level: {random resize and crop, random horizontal flip, Gaussian blur}, and the combination of text-level and image-level augmentations.
Below is the detailed descriptions of them:
Text-level:
  • Synonym Replacement: Following Bhattacharjee et al. (2023), only words identified as nouns, adjectives, adverbs and verbs will be considered for synonym replacement. We randomly choose $n$ words of these parts-of-speech to be replaced, where $n=10\%\times$ the number of words in the sentence. We use nltk.corpus.wordnet to obtain the synonyms and randomly choose one for each original word.
  • Random Swap: We randomly swap $n$ pair of words in the sentence, where $n=10\%\times$ the number of words in the sentence. This is achieved by using the nlpaug package Ma (2019).
  • Random Crop: We randomly remove $n$ consecutive words from the sentence, where $n=10\%\times$ the number of words in the sentence. This is achieved by using the nlpaug package Ma (2019).
Image-level:
  • Random Resize and Crop: We crop a random region of the image resize it to (224, 224).
  • Random Horizontal Flip: We flip the image horizontally with a probability of 0.5.
  • Gaussian Blur: We blur the image with the random Gaussian kernel. The kernel size range is from 5 to 9. The standard deviation range is from 0.1 to 5.0. All of the image-level augmentations are achieved by using torchvision.transforms.
Our experiments (Table 4) show that Gaussian blur results in the best overall performance. Thus, in our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the positive sample of the anchor item.
Table 4: Evaluation on different augmentation strategies of ConDA-TTA on Twitter-COMMs. Performances are reported in accuracy. The best performance is in bold text.
Augmentation Target Domain
M Cl Cv Avg
Synonym Replacement 79.7 80.2 80.3 80.07
Random Swap 79.9 80.2 80.3 80.00
Random Crop 79.3 80.5 80.0 79.93
Random Resize and Crop 80.1 80.2 79.9 80.07
Random Horizontal Flip 79.4 80.7 80.0 80.03
Gaussian Blur 80.0 80.4 80.1 80.16
Synonym Replacement 79.5 80.4 80.5 80.13
+ Gaussian Blur
Multimodal Feature Encoder Selection: We experiment with (i) BLIP-2 (Li et al., 2023c) multimodal and (ii) unimodal feature extractor, (iii) CLIP unimodal feature extractor and (iv) ALBEF (Li et al., 2021a) multimodal feature extractor. For (ii) and (iii), we experiment with vector concatenation and element-wise multiplication to obtain the multimodal feature.
We then train a one-layer linear classifier with the four features as inputs respectively and evaluate the performances (Table 5). Experimental results show that BLIP-2 multimodal features gives the best performance (in terms of accuracy) on both the training and test sets, as well as having the significant domain adaptation gap. We also evaluate the zero-shot performance of the BLIP-2 Image Text Matching head which only outputs a similarity score rather than the feature vector. We find it underperforms (achieving around 0.5 in accuracy on both training and test set) using BLIP-2 multimodal feature directly.
Table 5: Evaluation on different multimodal feature encoders of ConDA-TTA on Twitter-COMMs. Performances are reported in accuracy.
Multimodal Feature Encoder Target Domain
M Cl Cv
BLIP-2multi (supervised) 80.7 81.5 81.6
BLIP-2multi (source only) 72.4 78.8 77.9
$\Delta$ domain adaptation gap 8.3 2.7 3.7
ALBEF (supervised) 69.3 66.8 67.3
ALBEF (source only) 63.7 64.3 66.0
$\Delta$ domain adaptation gap 5.6 2.5 1.3
CLIP (supervised) 76.6 79.3 78.6
CLIP (source only) 72.1 77.3 78.1
$\Delta$ domain adaptation gap 4.5 2.0 0.5
BLIP-2uni (supervised) 72.5 75.0 74.4
BLIP-2uni (source only) 58.8 71.3 72.1
$\Delta$ domain adaptation gap 13.7 3.7 2.3
A.2 Computing Infrastructure
Our experiments are conducted on one NVIDIA A40 GPU. Each epoch takes around 20 mins to finish.
A.3 Sensitivity Analysis
The contrastive loss, MMD and TTA are three crucial components in our approach. This section briefly studies how the weights, $\lambda_{ctr}$ and $\lambda_{MMD}$, in the loss function and the batch size $b$ in TTA affect the performance of ConDA-TTA.
Figure 4: ConDA-TTA’s performances (in Acc) with different $\lambda_{MMD}$ and $\lambda_{ctr}$ values. The legend shows the target domain.
Figure 5: ConDA-TTA’s performances (in Acc) with different batch sizes. The legend shows the target domain.
Figure 4 illustrates ConDA-TTA’s performances with $\lambda_{MMD}=\{0.5,1,2,5\}$ and $\lambda_{ctr}=\{0.1,0.5,1,2\}$. When setting $\lambda_{MMD}$ to 1, our model achieves the best performance in 4 (M, Cl, B, U) out of 7 domain adaptation settings, and achieves close to best performance in the remaining 3 settings. Overall, $\lambda_{MMD}$ being too high causes the accuracy to drop a little. This might be because the model has put relatively less emphasis on the cross-entropy losses in Eq. (9). For $\lambda_{ctr}$, we find that setting it to 0.5 achieves the best performance in 5 settings. we find that when its value goes up, the model performance tends to decrease gradually. We conjecture this is because the distances (Section 4.2) between news with the same label have become much larger, making it more difficult to classify.
Figure 5 illustrates ConDA-TTA’s performances with test set’s batch size $b=\{64,128,256,512\}$. Overall, if computing resources permit, larger batch size tend to result in better model performances on most domain adaptation settings. However, the differences are not much. We argue that larger batch size helps the TTA to learn a less biased target domain statistics, because it includes more data points in one batch. However, we also conclude that it’s influence is limited.